Please complete all questions. Question 5: Let us say that the U.S. pays import subsidy to import sunflower seeds. Let the free market price of sunflower se
Answer #1

The free market price of $10 is shown in both graphs. At free market price of $10, quantity produced= 10 million pounds and quantity imported = 20 million pounds. Total quantity consumed thus 20+25= 45 million pounds.

As import subsidy increases the price to $11.50 (shown in market of partner country), the consumers in US pay only $11.50 - $2 (import subsidy)= $9.50 (shown in US market).

Under import subsisdy price of $9.50, quantity produced= 10 million pounds and quantity imported = 50 million pounds. Total quantity consumed thus 10+ 50=60 million pounds.

(a) US Partner Country $11.501 $100 a $9.50 10 25 4560 (6).(i) change in consumer surplus in us = Area a tbtc = (0.5)(0.50) (